### v1

# Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures

[…]
figurec
Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures
Oliver J. Clark oliver.clark1@monash.edu School of Physics and Astronomy, Monash University, Clayton, VIC, Australia    Anugrah Azhar Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PY, UK Physics Study Program, Faculty of Science and Technology, Syarif Hidayatullah State Islamic University Jakarta, Tangerang Selatan 15412, Indonesia    Ben A. Chambers Flinders Microscopy and Microanalysis, Flinders University, Bedford Park, South Australia 5042, Australia    Daniel McEwen School of Physics and Astronomy, Monash University, Clayton, VIC, Australia ARC Centre for Future Low Energy Electronics Technologies, Monash University, Clayton, VIC, Australia    Thi-Hai-Yen Vu School of Physics and Astronomy, Monash University, Clayton, VIC, Australia    M. Tofajjol H. Bhuiyan School of Physics and Astronomy, Monash University, Clayton, VIC, Australia    Rodion V. Belosludov Institute for Materials Research, Tohoku University, Sendai, 980-8577, Japan    Aaron Bostwick Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA    Chris Jozwiak Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA    Eli Rotenberg Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA    Seng Huat Lee Department of Physics, Pennsylvania State University, University Park, PA, 16802, USA 8 2D Crystal Consortium, Materials Research Institute, Pennsylvania State University, University Park, PA, 16802, USA    Zhiqiang Mao Department of Physics, Pennsylvania State University, University Park, PA, 16802, USA 8 2D Crystal Consortium, Materials Research Institute, Pennsylvania State University, University Park, PA, 16802, USA    Geetha Balakrishnan Department of Physics, University of Warwick, Coventry CV4 7AL, United Kingdom    Federico Mazzola Department of Molecular Sciences and Nanosystems, Ca Foscari University of Venice, Venice IT-30172, Italy CNR-SPIN UOS Napoli, Complesso Universitario di Monte Sant’Angelo, Via Cinthia 80126, Napoli, Italy    Sarah L. Harmer Institute for Nanoscale Science and Technology, Flinders University, Bedford Park, South Australia 5042, Australia Flinders Microscopy and Microanalysis, Flinders University, Bedford Park, South Australia 5042, Australia    Michael S. Fuhrer School of Physics and Astronomy, Monash University, Clayton, VIC, Australia ARC Centre for Future Low Energy Electronics Technologies, Monash University, Clayton, VIC, Australia    M. Saeed Bahramy Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PY, UK    Mark T. Edmonds School of Physics and Astronomy, Monash University, Clayton, VIC, Australia ARC Centre for Future Low Energy Electronics Technologies, Monash University, Clayton, VIC, Australia ANFF-VIC Technology Fellow, Melbourne Centre for Nanofabrication, Victorian Node of the Australian National Fabrication Facility, Clayton, VIC 3168, Australia
Van der Waals materials enable the construction of atomically sharp interfaces between compounds with distinct crystal and electronic properties Pham et al. (2022). This is dramatically exploited in moiré systems, where a lattice mismatch or twist between monolayers generates an emergent in-plane periodicity, giving rise to electronic properties absent in the constituent materials Devakul et al. (2021); Guo et al. (2023). In contrast, vertical superlattices—formed by stacking dissimilar van der Waals materials in the out-of-plane direction on the nanometer scale—remain largely unexplored, despite their potential to realize analogous emergent phenomena in three dimensions Wan et al. (2024). Here, we employ angle-resolved photoemission spectroscopy and density functional theory to investigate six-to-eight layer transition metal dichalcogenide (TMD) heterostructures. We find that even a single superlattice unit, i.e. a pair of stacked few-layer materials, exhibits emergent bands that closely resemble the hallmark three-dimensional electronic structure of many-layer TMDs, while simultaneously exhibiting an orbital- and momentum-dependent delocalization through the entire heterostructure due to its asymmetric architecture. We demonstrate how non-trivial band topology and unique TMD fermiologies are readily stabilized through these mechanisms, thus providing direct evidence for the novel electronic states possible through vertical superlattice engineering. These findings establish asymmetric heterostructures as platforms to invoke and control unprecedented combinations of the diverse quantum phases native to many-layer TMDs.
Figure 1: Chemical nature of an anisotropic transition metal heterostructure. (a) Schematic representation of a (WSe2)3/(MoSe2)3 heterostructure. Due to the $D_{3h}$ symmetry of each monolayer, a net dipole moment ±$\pm\mathbf{P}$ is induced, oriented perpendicular to the stacking axis, and oppositely aligned with respect to the adjacent layer. (b) and (c) Out-of-plane spin-projected electronic band structures for monolayers of WSe2 and MoSe2, respectively. The parameter $\Delta E_{\text{SO}}^{\text{W/Mo}}$ denotes the valley-dependent spin-orbit splitting at the K-point of the Brillouin zone. (d) and (e) Photoemission spectra collected from isolated 3ML thick flakes of WSe2 and MoSe2 (using 108 and 64 eV photons, respectively, and symmetrized about $\Gamma$) .The bands at $\Gamma$ are numbered from low energy to high. (f) and (g) Schematic energy diagrams illustrating the bonding and anti-bonding interactions at the K and $\Gamma$ points, respectively, with their orbital character predominantly influenced by the in-plane $d_{||}=\{{d_{xy},d_{x^{2}-y^{2}}}\}$ and out-of-plane $d_{z^{2}}$ orbitals of W/Mo. (h) Orbital-projected electronic structure of the (WSe2)3/(MoSe2)3 heterostructure, highlighting the distinct contributions from the constituent layers. The bands at $\Gamma$ are numbered from low energy to high.
Transition metal dichalcogenides (TMDs) are popular components for functional devices due to the ease in which their vast array of intrinsic properties can be augmented Chhowalla et al. (2013); Ahn (2020); Ciarrocchi et al. (2022); Manzeli et al. (2017); Regan et al. (2022). This is facilitated through anisotropic bonding, enabling both the isolation of TMD flakes of arbitrary thicknesses and their subsequent recombination into stacks with atomically sharp interfaces without the need for epitaxial registry Lisi et al. (2021); Pei et al. (2022); Nunn et al. (2023); Stansbury et al. ; Jones et al. (2022). These procedures can both alter the electronic properties inherent to the component compounds, and induce otherwise-inaccessible correlated phenomena through the fine tuning of the laboratory-defined relative orientations, or ‘twist angle’ Devakul et al. (2021); Guo et al. (2023). Despite van der Waals bonding, however, the interlayer hopping along the $c$-axis in TMDs is significant, driving orbital-dependent bandwidths of up to several electron volts in bulk systems Clark et al. (2018, 2022); Riley et al. (2014); Watson et al. (2024). This is yet to be exploited in artificial heterostructures, with current research largely rooted to homo- and heterobilayer systems, and thus to TMD properties in the 2D limit. This precludes, for example, the fine-tuning of band gaps through band quantisation Splendiani et al. (2010); Mak et al. (2010); Sun et al. (2016), the induction of $k_{z}$-mediated topological phases Bahramy et al. (2018), or the modification of charge density wave and magnetic transitions dependent on interlayer coupling Zhang et al. (2025); Watson et al. (2019); Feng et al. (2018).
Here, through nano-focused angle-resolved photoemission (nano-ARPES) and first principles calculations, we showcase vertical TMD heterostructures with emergent three-dimensional electronic band structures. Specifically, our 6-8 layer heterostructures, constructed simply by interfacing pairs of few-layer WSe2, MoSe2 and NbSe2 flakes, possess an orbital-dependent band localisation along the asymmetric $c$-axis. In this scenario, three-dimensional bands delocalised throughout the heterostructure coexist and interplay with layer-locked two-dimensional states to produce tunable arrays of material properties unlike those in either natural systems or artificial heterobilayers. We give special focus to $k_{z}$-mediated topological phase transitions, triggered by the interfacing of the few-layer TMD flakes, which produce pairs of distinct topologically non-trivial surfaces that are each defined largely by a single component material, unlike any naturally existing topological system. Similar phenomena should be expected for asymmetric heterostructures constructed from any pair of TMDs, and thus our findings unveil an enormous parameter space within which customizable three-dimensional electronic frameworks, mixing 2- and 3D properties from across the TMD family, can be produced and fine-tuned.
[…]
Figure 1 overviews the mechanism underpinning our findings. Fig.1(a) illustrates the lattice structure of a 6-layer MoSe2-WSe2 heterostructure, composed of three WSe2 layers stacked atop three MoSe2 layers along the crystalline $c$-axis. Each monolayer possesses $D_{3h}$ symmetry, with an in-plane polarisation $\bf{P}$ arising from M–Se dipoles (where M = Mo or W). The 2H stacking sequence results in a $2\pi$ rotation of adjacent layers, aligning their polar axes anti-parallel. In this heterostructure, the polarisation $\bf{P}$ of MoSe2 layers slightly exceeds that of WSe2 layers, reflecting the greater electronegativity difference between Mo and Se compared to W and Se. However, the stronger atomic spin-orbit interaction (SOI) of W dominates the spin physics of the system. Coupled with $\bf{P}$, this SOI generates an out-of-plane spin-orbit field, $B_{\text{SO}}\propto(\bf{P}\times\bf{k})_{z}$, which induces a Zeeman-like spin splitting of energy bands. The effect is most pronounced at the K points of the Brillouin zone, as shown in Fig. 1(b) and (c), where the spin splitting in WSe2 ($\Delta E^{\text{W}}_{\text{SO}}$) is more than twice that in MoSe2 ($\Delta E^{\text{Mo}}_{\text{SO}}$).
At the interface, the mismatch in spin-orbit coupling strengths gives rise to a hierarchical valley-spin polarisation. The MoSe2 valence bands are entirely contained within the $\Delta E^{\text{W}}_{\text{SO}}$ gap of WSe2, as shown in Fig. 1(f). This behaviour reflects the in-plane orbital nature of the K-point states, which are dominated by $d_{\parallel}={d_{xy},d_{x^{2}-y^{2}}}$ orbitals. These orbitals exhibit negligible out-of-plane coupling, suppressing interlayer hybridisation and preserving the monolayer-like electronic structure at K. By contrast, states at the $\Gamma$ point are dominated by $d_{z^{2}}$ orbitals, which extend into the van der Waals gap and exhibit substantial interlayer overlap. This leads to strong hybridisation at $\Gamma$, forming a broad bonding–anti-bonding energy window across the heterostructure, as depicted in Fig.1(g). While time-reversal symmetry suppresses spin splitting at $\Gamma$, the resulting band structure at this point is highly sensitive to interlayer coupling, in contrast to the robust monolayer-like dispersion retained at K. This dichotomy is evident in the full low-energy band structure of the 6-layer system (Fig.1(h)), where the K valleys retain their original characteristics while the $\Gamma$ valley forms a cascade of hybridised bands.
This stark contrast in orbital-dependent coupling is a hallmark of 2H-type transition metal dichalcogenides generally. The details of the $d_{\parallel}$ K-point valence bands are largely equivalent in monolayer and bulk 2H-TMD systems due to their minimal out-of-plane hopping potential Riley et al. (2015); Mak et al. (2010); Xiao et al. (2012), whereas $d_{z^{2}}$-derived $\Gamma$-point states exhibit substantial out-of-plane coupling to drive significant $k_{z}$ dispersion in bulk materials  Riley et al. (2014); Latzke et al. (2015); Kim et al. (2016); Alarab et al. (2023); Le Févre et al. (2024); Chang et al. (2014), thus underpinning the indirect-to-direct bandgap transition observed in few-layer systems Mak et al. (2010); Sun et al. (2016); Ernandes et al. (2021). Indeed, even isolated 3ML TMDs (Fig. 1(d-e)) have severely-altered energetics relative to their monolayer variants, but only around $\Gamma$.
This disparity highlights the importance of $\Gamma$-derived orbitals in band engineering, where artificial heterostructures can exploit interlayer hybridisation to tune the out-of-plane ($\parallel k_{z}$) dispersion by controlling layer composition, ordering, and alignment. Indeed, in the above simple case of a 6-layer MoSe2-WSe2 heterostructure, though discretized, the bonding-antibonding formation of states at and around the $\Gamma$ point closely approximate a $k_{z}$-integrated bulk band structure of a bulk 2H-TMD system. This implies that, with only a few layers, the bulk band dispersion can be significantly restored and even controllably altered to exhibit a mixed characteristic between two or more materials. Access to an extra dimension within which electronic properties can be tailored and controlled is of great relevance for nano-quantum device applications, where miniaturisation plays a major role. More obvious approaches to bulk-like band engineering would rely on the assembly of numerous monolayers with fixed relative orientations, stacked sequentially and periodically to replicate bulk periodicity in momentum space through the curation of a well-defined unit cell. By bypassing this labor-intensive approach to instead employ much simpler, two-component heterostructures in the intermediate layer regime, the same essential three-dimensional electronic behavior can be accessed within an architecture that lends itself more naturally to fine tuning with e.g. layer thickness or twist angles.
To demonstrate how this concept of achieving bulk-like band characteristics with mixed material properties can be realised in real systems, Figure 2 overviews the electronic structure of a (WSe2)3/(MoSe2)5 (Fig. 2(a)) and a (MoSe2)3/(WSe2)3 (Fig. 2(e)) heterostructure constructed by interfacing two few-layer TMD flakes on graphite/n-type silicon substrates (see Methods). This pair of heterostructures thus closely approximate opposite surfaces of a single MoSe2-WSe2 system. Photoemission spectra acquired from these heterostructure regions are displayed in Fig. 2(b, f).
As exemplified in Fig. 1(b-e), there is a one-to-one correspondence between the number of discrete $k_{z}$ sub-bands within the $d_{z^{2}}$-manifold at the $\Gamma$ point of 2H-structured TMDs and the number of monolayers within the system, Chang et al. (2014); Le Févre et al. (2024). However, as basic band theory stipulates that periodicity in momentum space cannot be achieved without periodicity in real space, the number of discrete $k_{z}$ sub-bands visible within the stacked regions would not be expected to reflect the full extent of the heterostructures which lack a well-defined unit cell. Yet, the photoemission spectra in Fig. 2(b,f) demonstrate that the number of sub-bands is not only increased relative to equivalent datasets acquired from the constituent flakes (shown in full in Supplemental Fig. 1), but also matches the total layer number in real space (i.e. eight for (b) and six for (f)) of the entire heterostructure. The electron mean free path of less than 0.7 nm, applicable to our photoemission experiments, precludes that the experimental result originates from a superposition of spectra from weakly-interacting top and bottom TMD flakes which each have thicknesses of $\approx$ 0.7 nm per monolayer Pan and Zahn (2022). This alternate scenario is further discarded through constant-energy $k_{x}$-$k_{y}$ contour mapping in Supplemental Figure 2 sup . Instead, the six or eight distinct $d_{z^{2}}$ sub-bands at $\Gamma$ are delocalised throughout the structure, and are therefore accessible to a surface sensitive probe.
While counterintuitive, these findings are entirely consistent with the discussion surrounding Figure 1: The array of hybridized bands at $\Gamma$ belonging to the collective band structure of the finite six or eight layer crystal structure closely mimics the discretised $k_{z}$-projected band manifold of ordered systems. For bands derived from orbitals with significant interlayer hopping potential, the $k_{z}$ dispersion is therefore effectively constructed without the need for a periodic repeating unit. To further test this description, the experimental band dispersions of both WSe2-MoSe2 heterostructures are compared to layer-projected, density functional theory (DFT)-based slab calculations in Fig. 2(c,g), (see Methods) with zero twist angle. The fine structure around the $\Gamma$ point is extremely well reproduced in each case, verifying the present interpretation while also demonstrating both that collective electron phases not captured by these calculations, and the large twist angles chosen experimentally in order to decouple Moiré effects (Sup. Fig. 3), are not conditional or consequential for the formation of these delocalised states.
To further scrutinize the observed three-dimensional behavior, in Figure 2(d) and (h) we track the evolution of the delocalised states at $\Gamma$ with changing photon energy, thus mapping their behavior with changing $k_{z}$ Damascelli (2004). These spectra can be compared to equivalent datasets for a three-layer MoSe2 and WSe2 flake (Sup. Fig. 1 (f,i)). For both the individual TMD flakes and the artificial heterostructures, each $d_{z^{2}}$-subband is two-dimensional (i.e. non-dispersive along $k_{z}$) in-line with their assignment as dimensionally confined states, though there is a photon energy dependence to their spectral weight. In the six and eight-layer hybrid systems (Fig. 2(d,h)), it becomes apparent that this shifting spectral weight between sub-bands is approximating the periodic $k_{z}$ dispersion that would be expected for a bulk 2H-structured TMD system. Indeed, the overlaid solid lines are a tight-binding model for an idealized bulk structure with alternating WSe2 and MoSe2 layers along the $c$-axis, once again showing how the periodic $k_{z}$-dispersions expected for ordered solids are emulated by interfacing pairs of TMDs in the intermediate layer regime.
[…]
The observed $c$-axis delocalisation, persisting even for this energetically-mismatched scenario, is facilitated by the lowering of the total energy of the system by partially lifting the constraints of 2D quantization. Along the $c$-axis, electrons form standing waves due to the finite thickness of the artificial crystals. This effect can be expected to vanish as one or both flakes approach the infinite-layer limit and electrons at the interface gain access to a bulk reservoir, ultimately preserving the quantized electronic structure of the remaining few-layer component. This is tested and shown explicitly in Supplemental Figure 5 through the characterization of a second overlap region within the sample in Fig. 2(i); between a 4ML-WSe2 flake atop a many-layer NbSe2 crystal. While charge transfer to the semiconducting top-layer still occurs in addition to some normalization of band features relative to the isolated 4ML-WSe2 top-layer, there are no additional well-defined $d_{z^{2}}$-subbands.
Figure 3: Spin-orbital variation in the (MoSe2)3/(WSe2)3 heterostructure. (a–c) Orbital- and spin-projected electronic structures calculated for both WSe2- and MoSe2-terminated configurations of the asymmetric (MoSe2)3/(WSe2)3 heterostructure. (d–f) Corresponding calculations for an idealised (MoSe2)3/(WSe2)3 heterostructure, where WSe2 and MoSe2 layers are alternately stacked along the $c$-axis.
While the presence of delocalised electron behavior within the $d_{z^{2}}$ orbital manifolds of the finite heterostructures is clear, the same is not true for regions of the electronic structure derived from in-plane orbitals. The inherent asymmetry to heterostructures constructed by interfacing two dissimilar TMD flakes ensures a spatial dependence to the broader electronic structure along the $c$-axis between limits defined by the distinct component materials. The scale of this dependence is dependent on orbital type. In Figure 3 and Supplemental Figs. 6-7, this real space evolution is explored in full. In Fig. 3(a), the orbitally-resolved WSe2 and MoSe2 terminations of the experimentally realized (MoSe2)3/(WSe2)3 heterostructure are compared. The valence band maxima at K and $\Gamma$, derived exclusively from $d_{\parallel}$ and $d_{z^{2}}$-orbitals respectively, represent the two extremes of $c$-axis delocalisation: The details of the spin-orbit splitting at K is determined only by the local crystal structure, while the band structure at $\Gamma$ is equivalent between terminations and indeed between the top and bottom layers (shown explicitly in Supplemental Figure 6). The rest of the band structure behaves similarly, with the real-space electronic structure dependence directly determined by the $d_{\parallel}$/ $d_{z^{2}}$ orbital ratio, further influenced by hybridization between bands of differing orbital origins.
The asymmetry of our artificial heterostructures can be expected to modify the spin structure relative to natural TMD systems. As discussed earlier, in a natural 2H-TMD, each constituent TMD layer, owing to its $D_{3h}$ symmetry, can only exhibit an out-of-plane spin-orbit field, whose direction is opposite to that of the neighbouring layers. Consequently, this implies that in-plane spin textures, including chiral spin-momentum locking types arising from the Rashba effect Bychkov and Rashba (1984), are effectively suppressed. Furthermore, the hidden out-of-plane spin textures of each layer can only become accessible at the surface, where translational symmetry breaking disrupts the natural cancellation of $B_{\text{SO}}$ along the stacking axis.
In contrast, our artificial structures can naturally accommodate both in-plane Rashba-type and out-of-plane Zeeman-type spin textures due to their unique heterostructural properties. Specifically, the chiral Rashba spin polarisation can emerge naturally from the potential gradient formed at the MoSe2/WSe2 interface and propagate throughout the entire structure via $d_{z^{2}}$ and other out-of-plane orbitals. To elucidate the implications of this, in Fig. 3(b-f), we compare a (MoSe2)3/(WSe2)3 heterostructure with an idealised six-layer (MoSe2/WSe2)3 system. The calculated layer-projected spin polarisation for the chiral ($x$) and out-of-plane ($z$) components is shown for both the asymmetric and idealised cases. As can be observed, in both configurations, strikingly similar patterns of in-plane and out-of-plane spin polarisation emerge, which remain valley-specific. Particularly, in both cases, the chiral ($x$) component is significantly enhanced around the $\Gamma$ point, where the $d_{z^{2}}$ contribution is dominant. As with the discussions surrounding the $k_{z}$ dispersions on show in Fig. 2, this again demonstrates how the key physics of ideally-stacked TMD heterostructures are very closely approximated by our simpler but asymmetric structures.
Figure 4: $k_{z}$-mediated topological phases in asymmetric heterostructures (a) Band dispersions along an axis close to K-$\Gamma$-K taken with 81 eV photons for (from left to right) bulk WSe2, 3ML-WSe2, a (WSe2)3/(MoSe2)5 heterostructure, 3ML MoSe2 and a (MoSe2)3/(WSe2)3 heterostructure. Dashed rectangles show the energy-momentum region covered by the plots (b-d.)(b) Zoom ins of the spectra in (a) with enhanced colour contrast. (c) Top-layer projected DFT calculations for the equivalent regions (d) Equivalent plots with resolved chiral ($\langle S_{x}\rangle$) spin component. Fully-formed (i.e. ungapped) topological surface states (TSSs) are indicated in (b) and (c).
Though orbital-dependent, the presence of three-dimensional electronic behavior in these types of artificial heterostructures prompts the question of whether surface-localised electronic states can exist where global crystalline symmetries are truncated. The disparate electronic structures of top and bottom surfaces would have profound implications for any surface states, should they exist. Fortunately, TMD heterostructures are particularly well-suited to explore this eventuality as the presence of surface states in bulk TMDs is enforced by delocalised electron behavior along $k_{z}$ Bahramy et al. (2018); Clark et al. (2018, 2019).
The families of three-fold symmetric 1T and 2H-structured TMDs possess a series of band inversions along their $k_{z}$ axes ($\parallel$ $\Gamma$-A) Huang et al. (2016); Yan et al. (2017); Noh et al. (2017); Li et al. (2017); Bahramy et al. (2018); Clark et al. (2018, 2019); Ghosh et al. (2019); Jiang et al. (2020); Mukherjee et al. (2020); Nurmamat et al. (2021); Clark et al. (2022); Chakraborty et al. (2023); Kimar et al. . These inversions, entirely contained to the chalcogen $p$-orbital manifold, generate up to five Dirac crossings (both bulk Dirac points and topological surface states (TSSs)) energetically stacked at the surface Brillouin centre. While the underpinning mechanism is equivalent for the dozen TMDs explicitly shown to host these states to date, the ladder of topological band crossings produced is delicately dependent on material-specific parameters, and thus are fingerprints of each individual system.
The minimal criteria to generate topological ladders in TMDs are only there exists a discrepancy in the $k_{z}$-bandwidths of in- and out-of-plane chalcogen $p$-orbital-derived states and $C_{3}$ symmetry is intact to maintain bulk Dirac crossings (though not crucial for the properties of surface Dirac states). The presence of bands well approximating a (discretised) $k_{z}$ dispersion along with orbital hopping asymmetries (discussed thus far for $d$-orbitals) imply that the above conditions are met in the experimental heterostructures considered here.
Figure 4 overviews the realisation of non-trivial band topology in our proof-of-principle MoSe2-WSe2 heterostructures. In 2H-TMDs like WSe2, there is one experimentally resolvable TSS centred $\approx$ 1.9 eV below the VBM Bahramy et al. (2018), as observed in spectra from the bulk region of flake used for device construction (Fig.  4(a)). This is compared to band dispersions from the pair of WSe2-MoSe2 heterostructures and their respective 3ML top layers. Experimental zoom-ins of the region wherein the bulk band inversion occurs and top-layer projected DFT calculations are provided in Fig. 4(b) and (c), respectively. While quasi-linear band dispersions are present in all cases, both 3ML TMD flakes undergo significant band renormalization when stacked atop a dissimilar flake. In 3ML-WSe2, both experiment and theory demonstrate a small local SOC-mediated gap between a pair of bands at the binding energy where the Dirac crossing occurs in the bulk limit. When interfaced to form a 3ML-WSe2/5ML-MoSe2 heterostructure, however, the lower of these two bands shifts to higher binding energy to give way to a gapless TSS with well-defined upper and lower branches which merge into the surrounding bands at higher $k_{\parallel}$, as for the case of bulk WSe2. Similarly contrasting behavior is evident for the second, thinner, heterostructure: While the smaller atomic SOC strength of Mo precludes resolving the finite band gaps for the isolated 3ML flake, a distinct crossing point emerges when stacked, differing from the states of 3L or 5ML-MoSe2 (Sup. Fig. 1). With reference to Supplemental Figs. 6 and 7, one can verify that the spectral weight of the Dirac points are in each case localised to the top and bottom monolayers only, demonstrating the surface states are a property of the collective lattice unit and not its individual components, in line with the $k_{z}$-mediated topological phase transition from which they derive Bahramy et al. (2018).
In contrast to $k_{z}$-mediated TSS found in naturally occurring TMDs, the appearances, energetic positions and characteristic chiral momentum-locked spin textures of these Dirac crossings (Fig. 4(d)), can differ between the two oppositely-orientated surfaces, with a $\approx 55$ meV discrepancy extracted for the specific (MoSe2)3/(WSe2)5 configuration from the layer-resolved calculations in Sup. Fig. 7, consistent with the experimental datasets from oppositely-orientated WSe2-MoSe2 structures in Fig. 4. This physics should apply to any pair of TMDs which are susceptible to $k_{z}$-mediated topological phase transitions, each with a unique set of Dirac crossings with surface-dependent behavior. For example, this could be utilized to tailor the five available Dirac cones in the 1T-structured TMD (semi)metals (Pd,Pt,Ni)(Se,Te)2, which occur near, or at, the Fermi level and have high sensitivity to relative hopping parameters in the unit cell Bahramy et al. (2018); Clark et al. (2019); Ferreira et al. (2021); Nicholson et al. (2021).
In conclusion, this work elucidates a new and straightforward method to engineer designer electronic structures in artificial bulk-like systems of transition metal dichalcogenides. By interfacing two flakes of few-layer TMDs, out-of-plane $k_{z}$-dispersions are generated between the limits of the parent compounds, facilitated by charge transfer where necessary. This produces a scenario in which the bulk-like band dispersions of 3D systems and the layer-locked electron behavior of 2D systems coexist within the same structure. These contrasting behaviors interplay, most notably in the production of $k_{z}$-mediated topological phase transitions that stabilize surface layer-localised topological states with properties set by the local crystal parameters, unlike any topologically non-trivial system existing in nature.
When combined with controlled lattice-mismatching, this approach could offer opportunities to engineer a coexistence of strong electronic correlations and non-trivial band topology that, while extremely rare naturally, are pathways towards exotic phases such as topological superconductivity, quantum spin liquids and topological Mott insulators Tokura (2022). Alternatively the non-equivalence of oppositely localized surface states, enforced by inherent $c$-axis structural asymmetry, could be exploited for 2D transistor-like components where the asymmetric surface electron transport can be controlled with band filling or layer-thickness. While demonstrated using flakes with sizes of only tens of microns, new methodologies for creating wafer-scale 2D material flakes Kang et al. (2017) could be exploited for this purpose. Together, the simplicity of this approach and the enormity of the phase space available together provide a new and accessible route towards a function-led construction of three-dimensional artificial systems.
Acknowledgements
O. J. C., M. S. F. and M. T. E. knowledge funding support from ARC Discovery Project DP200101345. M. T. E. acknowledges funding support from ARC Future Fellowship FT2201000290. O. J. C., D. M. and M.T.E. acknowledge travel funding provided by the International Synchrotron Access Program (ISAP) managed by the Australian Synchrotron, part of ANSTO, and funded by the Australian Government. D. M. was supported by the FLEET Centre of Excellence (ARC grant CE170100039). A.A., R.V.B. and M.S.B. gratefully acknowledge the Research Infrastructures at the Center for Computational Materials Science at the Institute for Materials Research for allocations on the MASAMUNE-IMR supercomputer system (Project No. 202112-SCKXX-0510) and MAHAMERU BRIN HPC facility under the National Research and Innovation Agency of Indonesia. M.S.B. acknowledges support from Leverhulme Trust (Grant No. RPG-2023-253). A.A. had support from the Indonesia Endowment Fund for Education (LPDP), NIB/202209223311735. R.V.B. and M.S.B. are grateful to E-IMR center at the Institute for Materials Research, Tohoku University, for continuous support. B. C. and S. L. H. acknowledge the facilities, and the scientific and technical assistance of Microscopy Australia (ROR: 042mm0k03) enabled by NCRIS and the Government of South Australia at Flinders Microscopy and Microanalysis (ROR: 04z91ja70), Flinders University (ROR: 01kpzv902). The work at the University of Warwick was supported by EPSRC, UK through Grant EP/T005963/1. This research used resources of the Advanced Light Source, which is a DOE Office of Science User Facility under contract no. DE-AC02-05CH11231. This work was performed in part at the Melbourne Centre for Nanofabrication (MCN) in the Victorian Node of the Australian National Fabrication Facility (ANFF)
Author Contributions OJC, DM, T-H-YV, MTHB, FM and MTE performed the nano-ARPES measurements. OJC and BC performed the nano-ESCA measurements. T-H-YV performed atomic force microscopy on the WSe2-NbSe2 heterostructure. AA, RVB and MSB performed the theoretical calculations. CJ, AB and ER provided user support at BL 7 MAESTRO of the Advanced Light Source. BC and SLH provided access for the nano-ESCA instrument at Flinders University, South Australia. SHL and ZM supplied semiconductors used in device construction GB supplied the NbSe2 crystals used for device construction. OJC conceived the project and wrote the manuscript with significant contributions from MSB and MTE and further input from all other authors. OJC was responsible for the overall project planning and direction.
Data Availability
Data are available from the authors upon reasonable request. xxxx
I Supplemental Materials for: Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures
Identifying component TMD flakes for heterostructure construction
[…]
In Sup. Fig. 1(a-b), optical images of the component flakes of (WSe2)3/(MoSe2)5 (a) and (MoSe2)3/(WSe2)3 are presented. These flakes, exfoliated onto PDMS using a standard scotch tape method, have thicknesses determined through comparison to photoluminescence spectra and optical images of reference samples.
Heterostructures assembled from these flakes are constructed in such a way that regions of individual components remain exposed laterally (see e.g. Fig. 2(a) of the main text), allowing for photoemission spectra to be acquired directly under identical experimental conditions to those of the heterostructures. This procedure allows for a direct comparison of band structures between component flakes and stacked heterostructures.
In 2H-TMDs, there is a one-to-one correspondence between the number of distinct quantized $k_{z}$-subbands of the $\Gamma$-centered $d_{z^{2}}$ orbital manifolds and the total thickness in monolayers Splendiani et al. (2010). This provides direct in situ verification of flake thicknesses through the enumeration of these doubly-degenerate bands in ARPES spectra. In Sup. Fig. 1(c, d, g) band dispersions close to the $\Gamma$-K direction are displayed for 5ML-MoSe2, 3ML-MoSe2 and 3ML-WSe2, with 5, 3 and 3 quantized states, respectively, indicated explicitly with dashed lines in Sup. Fig. 1(c), (f) and (i), respectively.
The 3ML-MoSe2 and 3ML-WSe2 spectra are compared to density functional theory calculations (see Methods) along the $\Gamma$-K direction. Agreement is very strong across the energy-momentum range shown.
Finally, in Sup. Fig. 1 (f,i), photon energy dependent spectra are acquired for $k_{\parallel}=0$ for both 3ML-MoSe2 (f) and 3ML-WSe2 (i). The dashed line indicates the $k_{z}$ values where the corresponding in-plane spectra are taken in (d) and (g).
The $k_{z}$ dependent datasets shown here and in Fig. 2 of the main text are acquired by continuously varying the photon energy. The relationship between the photon energy, $h\nu$ and $k_{z}$ is determined using a free-electron final state approximation Damascelli (2004):
$k_{z}=\sqrt{\frac{2m_{e}}{\hbar^{2}}(V_{0}+E_{k}(\cos^{2}(\theta))},$
where the inner potential, $V_{0}$, is taken to be 13 eV for all $k_{z}$ spectra in this work, in line with the value found for photon energy dependent studies on numerous bulk 2H-TMD systems (e.g. Riley et al. (2014)).
The varying spectral weight between the individually non-dispersive $k_{z}$ sub-bands (labeled in (f) and (i)) approximates the bulk $k_{z}$ dispersion of the infinite layer system Le Févre et al. (2024). With increasing layer number, these quantized states form a continuous $k_{z}$ dispersion, and the in-plane band dispersions of the highly three-dimensional bands become significantly $k_{z}$-broadened, thereby appearing diffuse, reflecting the small electron mean free path perpendicular to the surface Damascelli (2004).
Lack of rotational disorder in asymmetric heterostructures
Figure S2: (a-b) K-$\Gamma$-K and K-M-K band dispersions as extracted from a three-dimensional ($k_{x}$, $k_{y}$, $E-E_{\text{F}}$) dataset of the eight-layer 3ML-WSe2/5ML-MoSe2 heterostructure. (c) Constant-energy $k_{x}$-$k_{y}$ contours at the energies indicated in (b). 81 eV photons were used for all panels
In Supplemental Figure 2, constant energy $k_{x}-k_{y}$ contours and extracted high symmetry $\Gamma$-K and K-M-K dispersions are shown for the (WSe2)3/(MoSe2)5 heterostructure. Despite the large relative twist angle in both heterostructures between the WSe2 and MoSe2 flakes (see Sup. Fig. 3 for methodology), there is no evidence for rotational disorder. Note in particular the absence of a second set of the spin-orbit split local $d_{xy}$ orbital-derived VBM from the K points. The observed splitting at K is determined by the local crystal structure from the top WSe2 layer, in line with the $d_{xy}$-orbital character and the discussions surrounding Figure 1 of the main text. There are no band features consistent with that of the lower MoSe2 flake. This precludes the scenario that the observed eight-layer spectrum in Sup. Fig. 2(a) and in Figs. 2 and 4 of the main text derive simply from a superposition of spectra from weakly-interacting top and bottom flakes. Indeed, the electron mean free path is on the order of 5.8  Å  for the 81 eV photons used here, comparable to the thickness of only a single 1H unit of a 2H-TMD. Instead all eight visible sub-bands are present in the experimentally accessible top layers. We note that the quality of the photoemission data in Fig. 2 and Sup. Figs. 1-5 reflect the high quality of our heterostructure devices and exfoliation methods (see Methods), though we note the presence of subtle image warping due to instrumentation limitations, explaining both the differing size of the K point valleys and a slight shadowing of the bands at high $k_{x}$ in Sup. Fig. 2(c).
Twist angle determination between top and bottom flakes
[…]
Supplemental Figure 3 shows three band dispersions from three different regions of the sample housing the (MoSe2)3/(WSe2)3 heterostructure discussed in the main text. Band dispersions and constant energy contours are displayed for three regions selected by moving the sample laterally without rotation: (MoSe2)3 (b,e), (WSe2)3 (c,f) and (MoSe2)3/(WSe2)3(d, g). Data are collected using a nano-ESCA instrument at Flinders University, South Australia, with 21.2 eV photons from a Helium plasma discharge lamp.
Despite the much lower photon energy and elevated measurement temperature (300 K), these spectra are entirely consistent with those presented in the main text. Note also the continued lack of rotational disorder (as discussed for the eight layer heterostructure in Sup. Fig. 2). Indicated in Sup. Fig. 3 (e-f) are the in-plane azimuthal rotations applied to the raw data to achieve the high symmetry alignment in the Figure. The twist angle can be determined simply by calculating the difference between the azimuthal angle for the top and bottom TMD flakes. The twist angle for this sample is large at 17.5 degrees, and therefore band folding associated with the Moiré potential should not be expected in this system. The twist angle for the eight layer WSe2-MoSe2 heterostructure is similarly large.
Supplementary datasets for (WSe2)4/(NbSe2)x
Figure S4: (a) K-$\Gamma$-K and K-M-K and $\Gamma$-M-$\Gamma$ band dispersions as extracted from a three-dimensional ($k_{x}$, $k_{y}$, $E-E_{\text{F}}$) dataset of the six-layer 4ML-WSe2/2ML-NbSe2 heterostructure. (c) Constant-energy $k_{x}$-$k_{y}$ contours at the energies indicated in (b). The Brillouin zone is overlaid in purple, and high symmetry points labeled. 91.5 eV photons were used for all panels.
Sup. Fig. 4 shows constant energy contour $k_{x}$-$k_{y}$ mapping and extracted high symmetry dispersions for the (WSe2)4/(NbSe2)2 region of the sample shown in Fig. 2(i) of the main text.
These data demonstrate how the Fermi pocket created by interfacing the WSe2 and NbSe2 flakes is present in both the first and second Brillouin zones. The constant energy contours also exhibit no rotational disorder and no K valleys consistent with isolated NbSe2, again demonstrating that all observed bands are present within the top layers.
[…]
Supplemental Figure 5 compares the three dimensional electronic structure of (WSe2)4/(NbSe2)2 as shown in Figure 2 of the main text, to equivalent datasets from (WSe2)4 and (WSe2)4/(NbSe2)x, $x>>2$. All spectra are referenced to the Fermi level.
The in-plane band dispersion (b, close to $\Gamma$-K) and $k_{z}$ dispersion (c, taken at $k_{\parallel}$=0) for 4ML-WSe2 is consistent with those found for 3ML-WSe2 in Supplemental Fig. 1, with one extra $k_{z}$ subband in the $d_{z^{2}}$ manifold, consistent with the increased layer number. These spectra are compared to the case where the same 4ML-WSe2 flake overlaps with a bulk NbSe2 crystal. The number of $k_{z}$-subbands is unchanged due to the bulk-nature of the bottom component, consistent with the arguments in the main text. Despite this, there are two notable changes relative to free-standing 4ML-WSe2. Firstly, there are additional band features, most notably the conical bands appearing between $k_{z}$-subbands 3 and 4. In Sup. Fig. 5d, an additional spectra taken with 70.5 eV photons is displayed to increase the spectral weight of this conical feature. The second major change is the shift of all bands to shallower binding energies, with the topmost $k_{z}$-subband forming a sole Fermi pocket. While tangential to the findings concerning intermediate-regime asymmetric heterostructures, this result suggests yet another potential approach towards band structure engineering of few-layer TMD flakes. With a careful choice of substrate, the fermiology and band structures can be significantly modified.
Sup. Fig. 5g-i again consider the (WSe2)4/(NbSe2)2 case discussed in the main text. An additional band dispersion using 70.5eV photons (g) is displayed for comparison to the bulk case (d), and shows how a similar conical feature is formed.
We note that, unlike for the WSe2-MoSe2 heterostructures, one can see that there is a superposition of the stacked region with the top 4ML-WSe2 flake here. This has two possible origins: the probing light spot is at the boundary of the overlap (purple) and WSe2 (red) areas, or the less-than-pristine overlap region, as seen in Sup. Fig. 5(a), precludes full and complete charge transfer between the flakes.
Regardless of the origin, the significant charge transfer in the overlap region ensures that the $d_{z^{2}}$-derived bands remain separated from those originating from the top layer alone, thus facilitating an unambiguous enumeration of bands belong to the stacked heterostructure. We also note that the presence of only one extra copy of the WSe2 K point valleys shows that the charge transfer in this region is binary, and thus the observed sub-bands cannot be explained by a series of superimposed band structures with slightly different chemical potentials.
Orbital-dependent $c-$axis localisation
In Supplemental Figures 6 and 7, full layer-resolved DFT calculations for the K-$\Gamma$-K axis is shown for the (MoSe2)3/(WSe2)3 (Sup. Fig. 6) and the (WSe2)3/(MoSe2)5 (Sup. Fig. 7) heterostructures. Consistent with the discussions in the main text and with reference to Fig. 3, bands deriving entirely from out-of-plane orbitals are unchanged through the heterostructure (though intensities do vary), whereas bands deriving entirely from in-plane orbitals are strongly layer-dependent.
The topological surface states (TSSs) discussed in the main text and shown in Fig. 4 are visible here only in the surface layers and not close to the interface between the two compound types. This demonstrates how the TSSs are a product of the quasi-$k_{z}$ dispersion of the full heterostructure, and do not arise from one or both of the components in isolation.
Figure S6: Layer-integrated and layer-projected density functional calculations for the (MoSe2)3/(WSe2)3 heterostructure
Figure S7: Layer-integrated and layer-projected density functional calculations for the (WSe2)3/(MoSe2)5 heterostructure